OpenStudy (anonymous):

a student makes a$8.75 purchase at the bookstore with a $20.00 bill. The store has no bills and gives the change in guarters and fifty-cent pieces. there are 30 coins in all. how many of each kind are there? (trying to find the euqation to this problem it has me stumped). I just don't get it

OpenStudy (whpalmer4):

The amount to be given in change is $20 - $8.75 = $11.75. There are 30 coins. Let Q be the number of quarters, and F be the number of fifty-cent pieces. We know two things: 25Q + 50F = 1175 (we'll do it in cents to be free of decimals) Q + F = 30 Solve by substitution or elimination.

OpenStudy (whpalmer4):

I hope the student has strong pockets :-)